OpenStudy (anonymous):

Mel can paint a house in m hours and Ava can paint a house in n hours. Together, the time in hours it takes to paint six equally sized houses is given by the complex rational expression: 6 over 1/m + 1/n simplify the complex fraction...

jimthompson5910 (jim_thompson5910):

You have the following \[\large \frac{6}{\frac{1}{m}+\frac{1}{n}}\] Multiply EVERY term by the inner LCD mn to clear out the inner fractions to get \[\large \frac{6mn}{n+m}\]

OpenStudy (anonymous):

very nice! thank you :)